# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ Appendix A Numerical methods for quantum diffusion calculations
In this appendix, we briefly describe the numerical methods used to calculate electronic structure and quantum transport properties for non-interacting electrons at zero frequency. In the framework of the linear response, the Kubo-Greenwood formula [30, 31] (A.1), allows charge carrier dc-conductivity to be calculated, taking into account the complexity of the electronic structure and all possible multiple scattering phenomena (for a review see, e.g., Refs. [28, 366, 365]). This approach can be used for calculations (A.2) in reciprocal space for periodic systems whose size is not too large, thus allowing diagonalization of the Hamiltonian, and (A.3) in real space with very large cell, where a static defect random distribution can be instilled using recursion-type methods.
@@ A.1.1 Quantum diffusion of charge carriers and dc-conductivity
In the Kubo-Greenwood approach to transport properties [30, 31, 32], quantum diffusion, i.e., the diffusivity $\mathcal{D}$, is related to the average square spreading, $\Delta X^{2}(E,t)$ of charge carriers at each energy $E$ and each time $t$ [58, 28]. The central quantities are the velocity correlation function, $C(E,t)$, of states of energy $E$ at time $t$,
@@
[…]
@@
The average spreading, $\Delta X$, is deduced from the average square spreading, $\Delta X=\sqrt{\Delta X^{2}}$. It is easy to show that $C(E,t)$ is related to quantum diffusion by the relation [32, 58, 28],
[…]
@@
where $e$ is the electron charge, $E_{F}$ the Fermi energy, $n$ the density of states, and $\mathcal{D}$ the diffusivity related to the average square spreading by the relation,
$\mathcal{D}(E_{F})=\frac{1}{2}\,\lim_{t\rightarrow\infty}\frac{{d}}{{d}t}\,\Delta X^{2}(E_{F},t)\approx\lim_{t\rightarrow\infty}\frac{\Delta X^{2}(E_{F},t)}{t}.$ (61)
It is sometimes useful to define the diffusion coefficient, $D(E,t)=\frac{1}{t}\Delta X^{2}(E,t)$ (Eq. (16)).
@@ A.1.2 Relaxation time approximation (RTA)
The effect of the inelastic scattering is treated in a phenomenological way. This relaxation time approximation (RTA) has been used successfully to compute [27, 38] conductivity in approximants of quasicrystals where quantum diffusion and localization effects play an essential role [45, 179] and conductivity in organic semiconductors [427]. Following previous works [58, 28, 27, 427, 373], we assume that the velocity correlation function $C_{i}(E,t)$ of the system with inelastic scattering is given by,
@@
where $C(E,t)$ is the velocity correlation of the system with elastic scattering (adatoms, vacancies, di-vacancies,…) but without inelastic scattering. Here the inelastic scattering time $\tau_{i}$ is the cutoff time of the weak localization effects also called dephasing time. As shown in Refs. [58, 28, 27, 427], the propagation given by this formalism is unaffected by inelastic scattering at short times ($t<\tau_{i}$) and diffusive at long times ($t>\tau_{i}$), as it must be. Using the $t=0$ initial conditions, $\Delta X^{2}(E,t=0)=0$ and $\frac{{d}}{{d}t}\Delta X^{2}(E,t=0)=0$, and performing two integrations by parts, we obtain from equations (59), (60), (61), and (62), [38]
@@
$L_{i}^{2}(E_{F},\tau_{i})$ $=$ $\frac{1}{\tau_{i}}\int_{0}^{\infty}\!\Delta X^{2}(E_{F},t)\,{\rm e}^{-t/\tau_{i}}\,\mathrm{d}t\,,$ (65)
where $L_{i}(E_{F},\tau_{i})$ is the inelastic mean free path and $\mathcal{D}(E_{F},\tau_{i})$ the diffusivity. $\Delta X^{2}(E,t)$ is calculated for the system with the Hamiltonian given in the main text which represents only elastic scattering due to static defects. The above equations treat the inelastic scattering in a way that is equivalent to the standard approximation in mesoscopic physics. Indeed, in the presence of inelastic scattering, it is usually assumed that, $L_{i}(E_{F})\approx\sqrt{\Delta X^{2}(E_{F},\tau_{i})}$, thus the conductivity is given by the Einstein formula with a diffusivity $\mathcal{D}(E_{F},\tau_{i})\approx\Delta X^{2}(E_{F},\tau_{i})/(2\tau_{i})$ [189], which is essentially equivalent to the above equations.
RTA deals with the effect of temperature on diffusivity (and thus on conductivity) through an inelastic scattering time $\tau_{i}$. In this calculation, the Fermi-Dirac distribution function is taken equal to its zero temperature value. This is valid provided that the electronic structure varies smoothly on the thermal energy scale $k_{\rm B}T$ and is non-zero at $E_{F}$ (metal case). We have checked this assumption in the case of graphene (See Supplemental Material of Ref. [38]). However for semiconductors, it is essential to take into account the thermodynamic average at finite temperature, as we did for the quantum transport calculation in monolayer TMDs (Sec. 21) and monolayer and multilayer black phophorene (Sec. 3.5.2). Therefore the conductivity and the mobility at room temperature can be calculated (see Sec. 3.5).
In 2D systems with defects, it is known from the localization theory [189] that the diffusivity always tends to zero at very large time $t$; therefore, in RTA, $\lim_{\tau_{i}\to\infty}\sigma(\tau_{i})=0$, and $\sigma$ reaches a maximum value for a finite $\tau_{i}$, i.e., for a finite $L_{i}$ from Eq. (65). Thus, at each energy, the microscopic diffusivity, $\mathcal{D}_{M}$, and microscopic conductivity, $\sigma_{M}$, are defined as the maximum value of $\mathcal{D}(\tau_{i})$ and $\sigma(\tau_{i})$, respectively,
@@
It is also interesting to have an estimate of the elastic mean free path $L_{e}$ and the inelastic mean free path $L_{iM}$ value corresponding to the diffusive regime, i.e., $\sigma(L_{iM})\approx\sigma_{M}$. We compute the elastic mean free path $L_{e}$ along the $x$-axis, from the usual phenomenological formula [38],
$L_{e}(E)=\frac{1}{V_{0}(E)}\,{\rm Max}_{\tau_{i}}\left\{\frac{L_{i}^{2}(E,\tau_{i})}{\tau_{i}}\right\}=\frac{2\mathcal{D}_{M}(E)}{V_{0}(E)},$ (67)
where the velocity $V_{0}$ is the slope of $L_{i}(\tau_{i})$ at very small $\tau_{i}$ (ballistic regime see Sec. 3.1). It is important to note that such a definition of $L_{e}$ is not very accurate, and this calculation can only give an order of magnitude of the average distance between two elastic scattering events. Indeed, the formula (67) is not always valid when the electronic structure is strongly modified by static defects. Moreover, $V_{0}$ is overestimated since the numerical calculations include not only the intraband terms but also the interband terms [28]. In the case of monolayer graphene, we have shown [29] that these latter increase $V_{0}$ by a factor of $\sqrt{2}$ which leads to an underestimation of the $L_{e}$. However, roughly speaking, $L_{e}$ is the $L_{i}$ value above which the conductivity curve $\sigma(L_{i})$ reaches the plateau of diffusive regime due to elastic scattering. To better define the $L_{i}$ values corresponding to the diffusive regime, we sometime define the lengths $L_{i1}$ and $L_{i2}$ such as: $\forall L_{i}\in[L_{i1};L_{i2}]$, $\sigma(L_{i})>0.9\sigma_{M}$. We also determine the value $L_{iM}$ such that $\sigma$ is a maximal, i.e.,  $\sigma(L_{iM})=\sigma_{M}$. For instance, the values of $L_{e}$, $L_{i1}$, $L_{iM}$, and $L_{i2}$ are shown in the Supplemental Material of Ref. [160] for different concentrations of defects in Bernal bilayer graphene (BLG). The results show that $L_{e}\leq L_{i1}$ with the same order of magnitude, and the ratio $L_{i2}/L_{i1}$ varies from 5-10 to very large values, depending on the type of defects and their concentrations.
A.1.3 Numerical calculations
@@
    (Sec. A.2) For crystals without defects (except the effect of inelastic scattering), the diagonalization of the Hamiltonian in reciprocal space can be done for cells containing up to several $10^{4}$ orbitals. That permits a direct calculation of $\Delta X(E,t)$. This approach allowed us to study the quantum diffusion properties of Bloch states, considering not only the Boltzmann terms but also the non-Boltzmann terms which we had shown to be important in the small velocity regime (Sec. 2.1). This approach, which requires the structure to be periodic, does not allow randomly distributed defects to be included directly in the Hamiltonian. The effect of scattering defects, whether elastic or inelastic, can only be studied within the RTA (A.1.2) by introducing a relaxation time $\tau$ (Eq. (74)) instead of $\tau_{i}$.
@@
    (Sec. A.3) With a random distribution of static defects, such as vacancies or adatoms, it is necessary to work in a very large supercell (at least several millions of orbitals per supercell), and diagonalization of the Hamiltonian, which includes elastic defects, is not possible. In this case, we use a recursion method in real space [33, 34, 35, 36, 37]. Inelastic defects, e.g., electron-phonon interactions, can then be treated by RTA with an inelastic relaxation time $\tau_{i}$ (A.1.2).
@@ A.2.1 Without defects
In pure crystals at zero temperature, once the band structure is calculated from the tight-binding Hamiltonian, the average quadratic spreading can be computed exactly in the basis of Bloch states [27, 28]. The average square spreading is the sum of two terms,
@@
$V_{\mathrm{B}}$ is the Boltzmann velocity in the $x$ direction. The semi-classical theory is equivalent to taking into account only this Boltzmann term. The second term (interband contributions), $\Delta X^{2}_{\mathrm{NB}}(E,t)$, is a non-ballistic (non-Boltzmann) contribution. It is due to the non-diagonal elements in the eigenstates basis $\{|n\mathbf{k}\rangle\}$ of the velocity operator $\hat{V}_{x}$ (Eq. (58)). From the definitions (57) and (58), we obtain [27, 28],
$\Delta X_{\mathrm{B}}^{2}(E,t)=\hbar^{2}\,t^{2}\left\langle\,\sum_{\mathbf{k}}\left|\langle n\mathbf{k}|\hat{V}_{x}|n\mathbf{k}\rangle\right|^{2}\right\rangle_{E_{n}({\mathbf{k})}=E}$ (70)
@@ A.2.2 Extent $L_{w}$ of the wave packets
The formula seen in the previous section can be used to estimate the extent $L_{w}$ of a wave packet discussed in the introduction to this article and in section 2.1. Roughly speaking, for a system without defects, $L_{w}$ at energy $E$ is the quadratic time average of the non-Boltzmann square spreading, $L_{w}(E,t)=\sqrt{\Delta X_{\mathrm{NB}}^{2}(E,t)}$ (Eq. (5)). In quasicrystals, quasiperiodic tilings and flat-band systems such as twisted bilayer graphene, for many energies (but not necessarily all energies), we have shown numerically that $\Delta X_{\mathrm{NB}}^{2}(E,t)$ saturates (or almost saturates) at large $t$ (see Fig. 2), resulting in a finite $L_{w}^{\infty}(E)$ value which is generally large with respect to the interatomic distance. This new length is characteristic of the defect-free structure, and should be essential when the Boltzmann velocity is low. Indeed, in the usual semi-classical approximation charge carriers are viewed as wave packets. Their minimal extent is $L_{w}^{\infty}(E)$ and their dynamics is ballistic between scattering events (elastic or inelastic) only if $L_{w}^{\infty}$ is smaller than the mean-free path due to defects (see condition (1) and Sec. 2.1).

### v2

@@ Appendix A Numerical methods for quantum diffusion calculations
In this appendix, we briefly describe the numerical methods used to calculate electronic structure and quantum transport properties for non-interacting electrons at zero frequency. In the framework of the linear response, the Kubo-Greenwood formula Kubo57 , Greenwood58 (A.1), allows charge carrier dc-conductivity to be calculated, taking into account the complexity of the electronic structure and all possible multiple scattering phenomena (for a review see, e.g., Refs. Mayou08 , Fan14 , Fan20 ). This approach can be used for calculations (A.2) in reciprocal space for periodic systems whose size is not too large, thus allowing diagonalization of the Hamiltonian, and (A.3) in real space with very large cell, where a static defect random distribution can be instilled using recursion-type methods.
@@ A.1.1 Quantum diffusion of charge carriers and dc-conductivity
In the Kubo-Greenwood approach to transport properties Kubo57 , Greenwood58 , Kubo91 , quantum diffusion, i.e., the diffusivity $\mathcal{D}$, is related to the average square spreading, $\Delta X^{2}(E,t)$ of charge carriers at each energy $E$ and each time $t$ Mayou00 , Mayou08 . The central quantities are the velocity correlation function, $C(E,t)$, of states of energy $E$ at time $t$,
@@
where $\langle\cdots\rangle_{E}$ is the average on states with energy $E$, ReA\,{A} is the real part of A{A}, $\hat{V}_{x}(t)$ and $\hat{X}(t)$ are the Heisenberg representation of the velocity operator $\hat{V}_{x}$ and the position operator $\hat{X}$ along $x$ direction at time $t$,
@@
The average spreading, $\Delta X$, is deduced from the average square spreading, $\Delta X=\sqrt{\Delta X^{2}}$. It is easy to show that $C(E,t)$ is related to quantum diffusion by the relation Kubo91 , Mayou00 , Mayou08 ,
$\frac{d}{{d}t}\left(\Delta X^{2}(E,t)\right)=\int_{0}^{t}C(E,t^{\prime})\,{d}t^{\prime}.$ (59)
@@
where $e$ is the electron charge, $E_{F}$ the Fermi energy, $n$ the total density of states (DOS) (spin up + spin down), and $\mathcal{D}$ the diffusivity related to the average square spreading by the relation,
$\mathcal{D}(E_{F})=\frac{1}{2}\,\lim_{t\rightarrow\infty}\frac{{d}}{{d}t}\,\Delta X^{2}(E_{F},t)=\frac{1}{2}\,\lim_{t\rightarrow\infty}\frac{\Delta X^{2}(E_{F},t)}{t}.$ (61)
It is sometimes useful to define a time-dependent diffusion coefficient,
$D(E,t)=\frac{1}{2t}\Delta X^{2}(E,t).$ (62)
The above formulas are correct for a time-independent Hamiltonian. Thus they can be used when the inelastic scattering is not taken into account, typically at zero temperature, and we have used them in two distinct cases. Firstly, when the complex structure may confine electronic states (in quasicrystal and twisted bilayer graphene) without defects. Secondly to account for static defects (vacancies or adsorbed atoms) that are thus included in the Hamiltonian and lead to a long-time diffusive regime with a finite diffusivity. On the other hand, we always treat inelastic scattering by the relaxation time approximation (RTA) explained in the next section.
@@ A.1.2 Relaxation time approximation (RTA)
In this approximation, the effect of inelastic scattering is treated in a phenomenological way, by introducing a inelastic relaxation time $\tau_{i}$, whereas elastic scatterers are included in the Hamiltonian. We have implemented the RTA in two different ways, numerically, which do not yield the same quantitative results, but which qualitatively produce similar behaviors. We now present these two approaches.
(i) The simplest way of translating the effect of inelastic scatering time is to consider the diffusion coefficients $D$ (Eq. (62)) to be constant for $t\geq\tau_{i}$, thus from (61),
$\mathcal{D}(E_{F},\tau_{i})\approx D(E_{F},t=\tau_{i})=\frac{1}{2\tau_{i}}\,\Delta X^{2}(E,t=\tau_{i}),$ (63)
where $\Delta X^{2}(E,t)$ can be numerically computed (Eq. (57)) with or without elastic scattering defects included in the Hamiltonian. The inelastic mean free path is then,
$L_{i}(E_{F})\approx\Delta X(E_{F},t=\tau_{i}).$ (64)
Although this approach is rather crude, as it does not take into account the effect of inelastic diffusion, the results are qualitatively correct. The above equations treat the inelastic scattering in a way that is equivalent to the standard approximation in mesoscopic physics Lee85 .
(ii) A better approach  Mayou00 , Mayou08 , Trambly06 , Ciuchi11 , Lacroix_these is to assume that the velocity correlation function $C_{i}(E,t)$ of the system with inelastic scattering is given by,
@@
where $C(E,t)$ is the velocity correlation of the system with elastic scattering (adatoms, vacancies, di-vacancies,…) but without inelastic scattering. Here the inelastic scattering time $\tau_{i}$ is the cutoff time of the weak localization effects also called dephasing time. As shown in Refs. Mayou00 , Mayou08 , Trambly06 , Ciuchi11 , the propagation given by this formalism is unaffected by inelastic scattering at short times ($t<\tau_{i}$) and diffusive at long times ($t>\tau_{i}$), as it must be. Using the $t=0$ initial conditions, $\Delta X^{2}(E,t=0)=0$ and $\frac{{d}}{{d}t}\Delta X^{2}(E,t=0)=0$, and performing two integrations by parts, we obtain from equations (59), (60), (61), and (65), Trambly13
@@
$L_{i}^{2}(E_{F},\tau_{i})$ $=$ $\frac{1}{\tau_{i}}\int_{0}^{\infty}\!\Delta X^{2}(E_{F},t)\,{\rm e}^{-t/\tau_{i}}\,dt\,$, (68)
where $L_{i}(E_{F},\tau_{i})$ is the inelastic mean free path and $\mathcal{D}(E_{F},\tau_{i})$ the diffusivity. $\Delta X^{2}(E,t)$ is calculated for the system with the Hamiltonian given in the main text which represents only elastic scattering due to static defects.
RTA deals with the effect of temperature on diffusivity (and thus on conductivity) through an inelastic scattering time $\tau_{i}$. In this calculation, the Fermi-Dirac distribution function is taken equal to its zero temperature value. This is valid provided that the electronic structure varies smoothly on the thermal energy scale $k_{\rm B}T$ and is non-zero at $E_{F}$ (metal case). We have checked this assumption in the case of graphene (See Supplemental Material of Ref. Trambly13 ). However for semiconductors, it is essential to take into account the thermodynamic average at finite temperature, as we did for the quantum transport calculation in monolayer TMDs (Sec. 3.5.1) and monolayer and multilayer black phophorene (Sec. 3.5.2). Therefore the conductivity and the mobility at room temperature can be calculated (see Sec. 3.5).
Note that as discussed in A.2.3 below, the relaxation time approximation can be applied to a combination of elastic scattering and inelastic scattering. Such a relaxation time approximation has been used successfully, e.g., to compute Trambly06 , Trambly13 conductivity in approximants of quasicrystals where quantum diffusion and localization effects play an essential role Berger93 , Trambly05 .
A.1.3 Microscopic conductivity $\sigma_{M}$ and elastic scattering length $L_{e}$
In 2D systems with static defects (resonant or non-resonant defects), it is known from the localization theory Lee85 that the diffusivity always tends to zero at very large time $t$; therefore, in RTA, $\lim_{\tau_{i}\to\infty}\sigma(\tau_{i})=0$, and $\sigma$ reaches a maximum value for a finite $\tau_{i}$, i.e., for a finite $L_{i}$ from Eq. (68). Thus, at each energy, the microscopic diffusivity, $\mathcal{D}_{M}$, and microscopic conductivity, $\sigma_{M}$, are defined as the maximum value of $\mathcal{D}(\tau_{i})$ and $\sigma(\tau_{i})$, respectively,
@@ A.1.3 Microscopic conductivity $\sigma_{M}$ and elastic scattering length $L_{e}$
It is also interesting to have an estimate of the elastic mean free path $L_{e}$ and the inelastic mean free path $L_{iM}$ value corresponding to the diffusive regime, i.e., $\sigma(L_{iM})\approx\sigma_{M}$. We compute the elastic mean free path $L_{e}$ along the $x$-axis, from the usual phenomenological formula by extending the short-time ballistic regime to microscopic diffusivity Trambly13 ,
$L_{e}(E)\approx\frac{1}{V_{0}(E)}\,{\rm Max}_{\tau_{i}}\left\{\frac{L_{i}^{2}(E,\tau_{i})}{\tau_{i}}\right\}=\frac{2\mathcal{D}_{M}(E)}{V_{0}(E)},$ (70)
where the velocity $V_{0}$ is the slope of $\Delta X(t)$ or $L_{i}(\tau_{i})$ at very small $t$ or $\tau_{i}$, respectively (i.e., in the ballistic regime, see Sec. 3.1). The elastic scattering time is thus,
$\tau_{e}(E)\approx\frac{L_{e}(E)}{V_{0}(E)}.$ (71)
It is important to note that such a definition of $L_{e}$ is not very accurate, and this calculation can only give an order of magnitude of the average distance between two elastic scattering events. Indeed, the formula (70) is not always valid when the electronic structure is strongly modified by static defects. Moreover, $V_{0}$ is overestimated since the numerical calculations include not only the intraband terms but also the interband terms Mayou08 . In the case of monolayer graphene, we have shown Trambly16 that these latter increase $V_{0}$ by a factor of $\sqrt{2}$ which leads to an underestimation of the $L_{e}$. However, roughly speaking, $L_{e}$ is the value of $L_{i}$ above which the conductivity curve $\sigma(L_{i})$ reaches the plateau of diffusive regime due to elastic scattering. To better define the values of $L_{i}$ corresponding to the diffusive regime, we sometime define the lengths $L_{i1}$ and $L_{i2}$ such as: $\forall L_{i}\in[L_{i1};L_{i2}]$, $\sigma(L_{i})>0.9\sigma_{M}$. We also determine the value $L_{iM}$ such that $\sigma$ is a maximal, i.e., $\sigma(L_{iM})=\sigma_{M}$. For instance, the values of $L_{e}$, $L_{i1}$, $L_{iM}$, and $L_{i2}$ are shown in the Supplemental Material of Ref. Jouda21 for different concentrations of defects in Bernal bilayer graphene (BLG). The results show that $L_{e}\leq L_{i1}$ with the same order of magnitude, and the ratio $L_{i2}/L_{i1}$ varies from 5-10 to very large values, depending on the type of defects and their concentrations.
A.1.4 Numerical calculations
@@
    (Sec. A.2) For crystals without defects (except the effect of inelastic scattering), the diagonalization of the Hamiltonian in reciprocal space can be done for cells containing up to several $10^{4}$ orbitals. That permits a direct calculation of $\Delta X^{2}(E,t)$. This approach allowed us to study the quantum diffusion properties of Bloch states, considering not only the Boltzmann terms but also the non-Boltzmann terms which we had shown to be important in the small velocity regime (Sec. 2.1). This approach, which requires the structure to be periodic, does not allow randomly distributed defects to be included directly in the Hamiltonian. The effect of scattering defects, whether elastic or inelastic, can only be studied within the RTA (A.1.2) by introducing a relaxation time $\tau$ (Eq. (78)) instead of $\tau_{i}$.
@@
    (Sec. A.3) With a random distribution of static defects, such as vacancies or adatoms, it is necessary to work in a very large supercell (at least several millions of orbitals per supercell), and diagonalization of the Hamiltonian, which includes elastic defects, is not possible. In this case, we use a recursion method in real space Mayou88 , Mayou95 , Roche97 , Roche99 , Triozon02 . Inelastic defects, e.g., electron-phonon interactions, can then be treated by RTA with an inelastic relaxation time $\tau_{i}$ (A.1.2).
@@ A.2.1 Without defects
In pure crystals at zero temperature, once the band structure is calculated from the tight-binding Hamiltonian, the average quadratic spreading can be computed exactly in the basis of Bloch states Trambly06 , Mayou08 . The average square spreading is the sum of two terms,
@@
$V_{\mathrm{B}}$ is the Boltzmann velocity in the $x$ direction. The semi-classical theory is equivalent to taking into account only this Boltzmann term. The second term (interband contributions), $\Delta X^{2}_{\mathrm{NB}}(E,t)$, is a non-ballistic (non-Boltzmann) contribution. It is due to the non-diagonal elements in the eigenstates basis $\{|n\mathbf{k}\rangle\}$ of the velocity operator $\hat{V}_{x}$ (Eq. (58)). From the definitions (57) and (58), we obtain Trambly06 , Mayou08 ,
$\Delta X_{\mathrm{B}}^{2}(E,t)=t^{2}\left\langle\,\sum_{\mathbf{k}}\left|\langle n\mathbf{k}|\hat{V}_{x}|n\mathbf{k}\rangle\right|^{2}\right\rangle_{E_{n}({\mathbf{k})}=E}$ (74)
@@ A.2.2 Extent $L_{w}$ of the wave packets
The formula seen in the previous section can be used to estimate the extent $L_{w}$ of a wave packet discussed in the introduction to this article and in section 2.1. Roughly speaking, for a system without defects, $L_{w}$ at energy $E$ is the quadratic time average of the non-Boltzmann square spreading, $L_{w}(E,t)=\sqrt{\Delta X_{\mathrm{NB}}^{2}(E,t)}$ (Eq. (5)). In quasicrystals, quasiperiodic tilings and flat-band systems such as twisted bilayer graphene, for many energies (but not necessarily all energies), we have shown numerically that $\Delta X_{\mathrm{NB}}^{2}(E,t)$ saturates (or almost saturates) at large $t$ (see Fig. 2), resulting in a finite $L_{w}^{\infty}(E)$ value which is generally large with respect to the interatomic distance. This new length is characteristic of the defect-free structure, and should be essential when the Boltzmann velocity is low. Indeed, in the usual semi-classical approximation charge carriers are viewed as wave packets. Their minimal extent is $L_{w}^{\infty}(E)$ and their dynamics is ballistic between scattering events (elastic or inelastic) only if $L_{w}^{\infty}$ is smaller than the mean free path due to defects (see condition (1) and Sec. 2.1).